Question 1-2 (Chap 18 High low method)

Sales salaries and commissions are $10,000 when 80,000 units are sold, and $14,000 when 120,000 units are sold. Using the high-low method:

Q1. What is the variable portion of sales salaries and commission?

Q2. What is the fixed portion of sales salaries and commission?
Answer #1

1.

Variable cost per unit = (Highest activity cost - Lowest activity cost)/(Highest activity - Lowest activity)

= (14,000 - 10,000)/(120,000 - 80,000)

= 4,000/40,000

= $0.10 per unit

variable portion of sales salaries and commission = $0.10 per unit

2.

Fixed cost = Highest activity cost - (Highest activity x Variable cost per unit)

= 14,000 - (120,000 x 0.10)

= 14,000 - 12,000

= $2,000

fixed portion of sales salaries and commission = $2,000